OpenStudy (anonymous):

What 5y+26=41 What is the y

OpenStudy (anonymous):

First to answer right gets the medal

OpenStudy (rottenywrm):

5y = 41 -26 y = 15/5 y = 3

OpenStudy (sbuck98):

if you need more help on problems I have the time!

OpenStudy (anonymous):

hehe lazy gideon

OpenStudy (astrophysics):

Seems someone already answered this for you, what's the problem?

OpenStudy (anonymous):

NOTHING

OpenStudy (astrophysics):

Haha alright, you bumped it so just making sure